# DynGMA: a robust approach for learning stochastic differential equations from data

### v1

@@ 5 Summary
In this paper, we present a novel method for approximating the transition density of a Stochastic Differential Equation (SDE) aimed at learning unknown stochastic dynamics from data. The proposed density approximation allows for multiple time steps, thereby obtaining a more accurate approximation for large time interval, while existing methods are usually based on one-step stochastic numerical schemes and thus necessitate data with sufficiently high time resolution. Our approach includes a Gaussian density approximation inspired directly by the random perturbation theory of dynamical systems blagoveshchenskii1962diffusion ; blagoveshchenskii1961certain , complemented by an extension named as dynamical Gaussian mixture approximation (DynGMA). The targeted SDE describes dynamical systems with random perturbations. While its drift and diffusion functions are fully unknown, discrete observations of certain trajectories are available. To reconstruct unknown stochastic dynamics, we leverage neural networks to approximate both the drift and diffusion functions of the SDE. Subsequently, we employ the proposed density approximation to construct a likelihood loss, which serves as the training objective for the neural networks. The proposed method is capable of reconstructing unknown governing functions of SDEs through a data-driven approach, and thus can be applied to invariant distribution computing from data. Benefiting from the robust density approximation, our method demonstrates robustness and possesses many benefits across various application scenarios. In particular, it exhibits superior accuracy with a given step size, remains effective even when confronted with trajectory data with low time resolution, and performs well even in the presence of higher levels of measurement noise. Finally, we validate the superior properties of our proposed method in comparison to competing baseline models in a variety of examples across various scenarios, including a biological network model and an epidemiological compartment model.
@@
There are many possible extensions of our work. As we have provided the error quantification of the Gaussian approximations, it is interesting to further explore the interplay between the training process and numerical integration to adapt the time step. A possible direction is to estimate the magnitude of the current diffusion and adaptively select the discretization accuracy based on it in the training process. Additionally, we would like to combine our density approximation with kernel methods hamzi2023note in future works. In terms of theory, the quantification of the error between the learned and target governing functions, as well as the analysis of convergence in subsequent invariant distribution computation, remains an open challenge. It is important to note that the error of the approximated transition density does not directly yield the aforementioned bounds. While there have been previous analyses on this task gu2023stationary , their analysis focuses on SDE with constant diffusion matrix, and the loss function is derived through the one-step Euler-Maruyama scheme. We would like to explore the feasibility of extending existing analysis results to our method in future research. In addition, it is worth mentioning that DynGMA operates independently of specific structures, and the $f_{\theta}$ and $\sigma_{\theta}$ can be any valid neural networks or functions, depending on the specific problem. Our algorithm naturally accommodates the encoding of partially known physical terms, such as generalized Onsager principle chen2023constructing ; yu2021onsagernet , symplectic structure bertalan2019learning ; greydanus2019hamiltonian , and GENERIC formalism zhang2022gfinns . Using our density approximation to learn structure-preserving stochastic dynamics might be an interesting direction. Moreover, in practical applications, many high-dimensional systems necessitate order reduction, requiring simultaneous learning of the latent states and the governing function. An example includes the learning of the macroscopic thermodynamic description of the stretching dynamics of polymer chains chen2023constructing . The investigation of the performance of our method in handling such tasks is also left for future work.
@@ Appendix B Computing invariant distributions from data
Our learning model is able to discover the dynamics and thus can be turned into an invariant distribution computation algorithm for the situation where we have access to noisy data but the governing equation is unknown. We next introduce the detailed algorithm following lin2023computing .
The density function $p(x)$ of the the invariant distribution of the SDE (1) satisfies the stationary FPK equation:
@@
Following lin2023computing , we define the generalized potential
@@
Then, by writing the the diffusion function as $\sigma(x)=\sqrt{2\varepsilon}\bar{\sigma}(x)$ where $\lVert\bar{\sigma}\rVert_{F}=1$ and $\bar{\Sigma}:=\bar{\sigma}\bar{\sigma}^{\top}$, $V(x)$ satisfies
@@
Focusing on constant diffusion (i.e., $\mathop{}\!\mathrm{d}\bar{\Sigma}\equiv 0$), lin2023computing proposes to identify the drift and diffusivity by parameterizing $f$ via (24a) and employing the Euler–Maruyama method to approximation transition density, and simultaneously compute the invariant distribution by network-based residual minimization according to (24b). Alternatively, we first identify a approximation of $f$ and $\bar{\Sigma}$, i.e., $f_{\theta^{*}}$ and $\sigma_{\theta^{*}}$ by optimizing (6) or (7) and then employ network-based residual minimization for solving (B). Specifically, we parameterize the generalized potential $V$ and the auxiliary function $g$ as $V_{\phi}$ and $g_{\phi}$ respectively. The parameters $\phi$ can be obtained by minimizing the physical-informed loss, i.e.,
@@
where $\varepsilon_{\theta^{*}}=\left\lVert\sigma_{\theta^{*}}\right\rVert_{2}^{2}/2$, $\bar{\Sigma}_{\theta^{*}}=\sigma_{\theta^{*}}\sigma_{\theta^{*}}^{\top}/2/\varepsilon_{\theta^{*}}$, and $y^{n}_{m}$ are the observations of given $N$ trajectories. Adam optimizer kinga2015method is applied to update the parameters. After training, we shift the potentials $V_{\phi^{*}}$ so that their minimum value in the domain is equal to zero.
Appendix C Proofs
@@
According to the facts that $\widetilde{\Sigma}(lh/L)=\mathcal{O}(\left\lVert\sigma_{\theta}\right\rVert_{F}^{2}h)$ and $\bar{\Sigma}_{l+1}$ is of order-2 approximation, we conclude the second estimate. Using the classical numerical analysis results (hairer1993solving, , Theorem 3.4, Page 160), the multi-step error reads

### v2

@@ 5 Summary
In this paper, we present a novel method for approximating the transition density of a Stochastic Differential Equation (SDE) aimed at learning unknown stochastic dynamics from data. The proposed density approximation allows for multiple time steps, thereby obtaining a more accurate approximation for large time interval, while existing methods are usually based on one-step stochastic numerical schemes and thus necessitate data with sufficiently high time resolution. Our approach includes a Gaussian density approximation inspired directly by the random perturbation theory of dynamical systems [4, 5], complemented by an extension named as dynamical Gaussian mixture approximation (DynGMA). The targeted SDE describes dynamical systems with random perturbations. While its drift and diffusion functions are fully unknown, discrete observations of certain trajectories are available. To reconstruct unknown stochastic dynamics, we leverage neural networks to approximate both the drift and diffusion functions of the SDE. Subsequently, we employ the proposed density approximation to construct a likelihood loss, which serves as the training objective for the neural networks. The proposed method is capable of reconstructing unknown governing functions of SDEs through a data-driven approach, and thus can be applied to invariant distribution computing from data. Benefiting from the robust density approximation, our method demonstrates robustness and possesses many benefits across various application scenarios. In particular, it exhibits superior accuracy with a given step size, remains effective even when confronted with trajectory data with low time resolution, and performs well even in the presence of higher levels of measurement noise. Finally, we validate the superior properties of our proposed method in comparison to competing baseline models in a variety of examples across various scenarios, including a biological network model and an epidemiological compartment model.
@@
There are many possible extensions of our work. As we have provided the error quantification of the Gaussian approximations, it is interesting to further explore the interplay between the training process and numerical integration to adapt the time step. A possible direction is to estimate the magnitude of the current diffusion and adaptively select the discretization accuracy based on it in the training process. Additionally, exploring the combination of DynGMA with kernel methods [18], as well as extending DynGMA to forward solving SDEs [27], could be interesting directions. In terms of theory, the quantification of the error between the learned and target governing functions, as well as the analysis of convergence in subsequent invariant distribution computation, remains an open challenge. It is important to note that the error of the approximated transition density does not directly yield the aforementioned bounds. While there have been previous analyses on this task [16], their analysis focuses on SDE with constant diffusion matrix, and the loss function is derived through the one-step Euler-Maruyama scheme. We would like to explore the feasibility of extending existing analysis results to our method in future research. In addition, DynGMA and its application in computing invariant distributions from data are specifically designed for unknown SDEs driven by Gaussian noise. We would like to derive methods capable of effectively handling general non-Gaussian noise in future research. It is worth mentioning that DynGMA operates independently of specific structures, and the $f_{\theta}$ and $\sigma_{\theta}$ can be any valid neural networks or functions, depending on the specific problem. Our algorithm naturally accommodates the encoding of partially known physical terms, such as generalized Onsager principle [11, 48], symplectic structure [3, 15], and GENERIC formalism [49]. Using our density approximation to learn structure-preserving stochastic dynamics might be an interesting direction. Moreover, in practical applications, many high-dimensional systems necessitate order reduction, requiring simultaneous learning of the latent states and the governing function. An example includes the learning of the macroscopic thermodynamic description of the stretching dynamics of polymer chains [11]. The investigation of the performance of our method in handling such tasks is also left for future work.
Acknowledgments
This project is supported by the National Research Foundation, Singapore, under its AI Singapore Programme (AISG Award No: AISG3-RP-2022-028) and the NRF fellowship (project No. NRF-NRFF13-2021-0005).
@@ Appendix B Computing invariant distributions from data
Our learning model is able to discover the dynamics and thus can be turned into an invariant distribution computation algorithm for the situation where we have access to noisy data but the governing equation is unknown. We next introduce the detailed algorithm following [29].
The density function $p(x)$ of the invariant distribution of the SDE (1) satisfies the stationary FPK equation:
@@
Following [29], we define the generalized potential
@@
Then, by writing the diffusion function as $\sigma(x)=\sqrt{2\varepsilon}\bar{\sigma}(x)$ where $\lVert\bar{\sigma}\rVert_{F}=1$ and $\bar{\Sigma}:=\bar{\sigma}\bar{\sigma}^{\top}$, $V(x)$ satisfies
@@
Focusing on constant diffusion (i.e., $\mathop{}\!\mathrm{d}\bar{\Sigma}\equiv 0$), [29] proposes to identify the drift and diffusivity by parameterizing $f$ via (24a) and employing the Euler–Maruyama method to approximation transition density, and simultaneously compute the invariant distribution by network-based residual minimization according to (24b). Alternatively, we first identify a approximation of $f$ and $\bar{\Sigma}$, i.e., $f_{\theta^{*}}$ and $\sigma_{\theta^{*}}$ by optimizing (6) or (7) and then employ network-based residual minimization for solving (24). Specifically, we parameterize the generalized potential $V$ and the auxiliary function $g$ as $V_{\phi}$ and $g_{\phi}$ respectively. The parameters $\phi$ can be obtained by minimizing the physical-informed loss, i.e.,
@@
where $\varepsilon_{\theta^{*}}=\left\lVert\sigma_{\theta^{*}}\right\rVert_{2}^{2}/2$, $\bar{\Sigma}_{\theta^{*}}=\sigma_{\theta^{*}}\sigma_{\theta^{*}}^{\top}/2/\varepsilon_{\theta^{*}}$, and $y^{n}_{m}$ are the observations of given $N$ trajectories. Adam optimizer [23] is applied to update the parameters. After training, we shift the potentials $V_{\phi^{*}}$ so that their minimum value in the domain is equal to zero.
Appendix C Learning Beneš SDE from data
Herein, we investigate the performance of DynGMA in learning Beneš SDE from data. As we focus on dynamical systems with random perturbations, we reduce the intensity of diffusion, considering
$\mathop{}\!\mathrm{d}x=\tanh x\mathop{}\!\mathrm{d}t+0.3\mathop{}\!\mathrm{d}\omega.$
We sample $5\times 10^{5}$ points from the uniform distribution on $[-1,1]$ as initial values and compute the states after time step $\Delta t$ to obtain the training dataset. We employ a feedforward neural network (FNN) with one hidden layer containing 64 units and sigmoid activation functions to parameterize the drift $f_{\theta}$. Additionally, we use a trainable constant to parameterize the diffusion $\sigma_{\theta}$. We use Algorithm 1 to approximate the transition density, setting the step size of each sub-Gaussian step to $h=0.2$ and the step size of discretization (14) to $0.1$. This configuration corresponds to $K=\lceil\Delta t/0.2\rceil$ and $L=\min(0.1,\Delta t)/0.1$. Specifically, for $\Delta t$ values of $0.1$, $0.2$, and $0.4$, the values of $K$ are $1,1,2$, and the values of $L$ are $1,2,2$, respectively. During training, we employed full-batch Adam optimization for $5\times 10^{3}$ epochs, with the learning rate set to exponentially decay with linearly decreasing powers from $0.1$ to $0.01$.
Settings $\Delta t=0.1$ $\Delta t=0.2$ $\Delta t=0.4$
$e_{\sigma}$ 2.37e-3$\pm$1.00e-3 7.44e-2$\pm$2.14e-3 8.78e-4$\pm$9.24e-4 2.37e-1$\pm$1.63e-1 2.42e-2$\pm$4.06e-3 3.30e-1$\pm$4.97e-3
$e_{f}$ 6.06e-3$\pm$2.03e-3 3.01e-2$\pm$2.43e-3 6.24e-3$\pm$1.37e-3 6.25e-2$\pm$1.08e-2 1.08e-2$\pm$3.32e-3 1.10e-1$\pm$1.16e-3
Table 4: Quantitative results for learning Beneš SDE. The results are obtained by taking the mean of $5$ independent experiments, and the errors are recorded in the form of mean $\pm$ standard deviation.
After the training process is completed, we calculate $e_{\sigma}$ and $e_{f}$ at $1000$ grid points within the interval $[-1,1]$, and document the errors in Table 4 to quantitatively evaluate accuracy. The findings demonstrate that DynGMA also achieves lower errors, consistent with the results presented in Section 4.
Appendix D Proofs
@@
According to the facts that $\widetilde{\Sigma}(lh/L)=\mathcal{O}(\left\lVert\sigma_{\theta}\right\rVert_{F}^{2}h)$ and $\bar{\Sigma}_{l+1}$ is of order-2 approximation, we conclude the second estimate. Using the classical numerical analysis results [17, Theorem 3.4, Page 160], the multi-step error reads